# When Are Bias-Free ReLU Networks Effectively Linear Networks?

### v2

@@ 5 Learning Dynamics in Deep Bias-Free ReLU Networks
In Section 4 we showed that two-layer bias-free ReLU networks behave like linear networks under symmetry Condition 3 and small initialization. This does not extend to deep bias-free networks. When trained on a dataset satisfying Condition 3, deep bias-free ReLU networks can learn nonlinear solutions if the target function is nonlinear, as shown in LABEL:fig:express-fan (upper right). Nonetheless, we find deep bias-free ReLU networks can form low-rank weights that are similar to those in deep linear networks. We give an example where the empirical input distribution is even and the target function is linear.
@@
where $u$ represents the norm of each layer, and ${\bm{r}},{\bm{r}}_{1},{\bm{r}}_{2},\cdots,{\bm{r}}_{L}$ are unit norm column vectors. The vectors ${\bm{r}}_{l}^{+},{\bm{r}}_{l}^{-}$ denote the positive and negative elements in ${\bm{r}}_{l}$. The equal norm $u$ of all layers is a consequence of small initialization (Du et al., 2018). Note that the weights can be written in blocks, as Equation 14, only after permuting the positive and negative elements. We use this permuted notation for the sake of exposition; no additional assumptions are required.
In a deep ReLU network, we empirically find that when the weights are trained from small initialization and the target function is linear, the weights form a particular rank-one and rank-two structure. The weights of a deep bias-free network can be written approximately as
@@
${\bm{W}}_{L}$ $=u{\bm{r}}_{L-1}^{\top}=u\begin{bmatrix}{{\bm{r}}_{L-1}^{+}}^{\top}&{{\bm{r}}_{L-1}^{-}}^{\top}\end{bmatrix}.$ (15c)
For the first and last layers, the weights in the deep ReLU network have the same rank-one structure as their linear counterpart. For the intermediate layers, weights in the deep ReLU network are rank-two. Specifically, positive weights in the ReLU network correspond to positive weights in the linear network and zero weights in the ReLU network correspond to negative weights in the linear network. We visualize the low-rank weights of a three-layer linear network and a three-layer bias-free ReLU network in Figure 4.
Figure 4: Low-rank weights in deep linear and ReLU bias-free networks. A three-layer linear network and a three-layer ReLU network are trained on the same dataset starting from the same small random weights. The dataset has a linear target function and an even empirical input data distribution. We plot the weights when the loss has approached zero. ${\bm{W}}_{1}$, ${\bm{W}}_{3}$, and positive elements in ${\bm{W}}_{2}$ have approximately the same structure in the linear and ReLU networks. Elements of ${\bm{W}}_{2}$ that are negative in the linear network are approximately zero in the ReLU network. The neurons are permuted for better visualization. Experimental details are provided in Appendix H.
If we assume $\|{\bm{r}}_{l}^{+}\|=\|{\bm{r}}_{l}^{-}\|,(l=1,2,\cdots,L-1)$, which is true when the network width approaches infinity, the deep bias-free ReLU network with weights defined in Equation 15 implements a linear function
@@
In the first equality, we drop the activation functions except the one between the first and second layers. This is because the second layer weights, ${\bm{W}}_{2}$, is non-negative as shown in LABEL:fig:deep-relu-rank12, and so is the output of a ReLU activation function, $\sigma({\bm{W}}_{1}{\bm{x}})$. Hence, their product, ${\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})$, is also non-negative. We thus have $\sigma({\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}))={\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})$. The same applies to all subsequent layers. The second equality is obtained by substituting the weights defined in Equation 15 into the expression.
When the empirical input distribution is even and the target function is linear, the learning dynamics of the deep bias-free ReLU network with weights as Equation 15 reduces to (see Appendix E)
$\tau\dot{\bm{W}}_{l}=\left(\prod_{l^{\prime}=l+1}^{L}{\bm{W}}_{l^{\prime}}\right)^{\top}\left(\frac{1}{2}{\bm{\beta}}^{\top}-\frac{1}{4}\prod_{l^{\prime}=1}^{L}{\bm{W}}_{l^{\prime}}{\bm{\Sigma}}\right)\left(\prod_{l^{\prime}=l-1}^{L}{\bm{W}}_{l^{\prime}}\right)^{\top}.$ (17)
Except for constant coefficients, these equations are the same as that of the deep linear network given in Equation 7. We show, in Appendix E, that weights which have formed a low-rank structure as defined in Equation 15 maintain the structure over training.
@@
We conjecture that in deep linear networks and certain deep ReLU networks, the weights of an intermediate layer align with the inputs to that layer. For example, the second-layer weight in a deep linear network is ${\bm{W}}_{2}^{\mathrm{lin}}=u{\bm{r}}_{2}{\bm{r}}_{1}^{\top}$ as given in Equation 14. Every row of ${\bm{W}}_{2}^{\mathrm{lin}}$ aligns with ${\bm{r}}_{1}^{\top}$, which is parallel to any input to the second layer, ${\bm{W}}_{1}^{\mathrm{lin}}{\bm{x}}=u{\bm{r}}_{1}{\bm{r}}^{\top}{\bm{x}}$. The second-layer weight in the deep ReLU network is given in Equation 15. Some rows of ${\bm{W}}_{2}$ align with $\begin{bmatrix}{{\bm{r}}_{1}^{+}}^{\top}&{\bm{0}}\end{bmatrix}$, which is parallel to some inputs (${\bm{r}}^{\top}{\bm{x}}>0$) to the second layer, $\sigma({\bm{W}}_{1}{\bm{x}})=u\begin{bmatrix}{\bm{r}}_{1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}$. Other rows of ${\bm{W}}_{2}$ align with $\begin{bmatrix}{\bm{0}}&{{\bm{r}}_{1}^{-}}^{\top}\end{bmatrix}$, which is parallel with inputs (${\bm{r}}^{\top}{\bm{x}}<0$) to the second layer, $\sigma({\bm{W}}_{1}{\bm{x}})=u\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}$. This alignment phenomenon has been proven for deep linear networks (Ji & Telgarsky, 2019; Marion & Chizat, 2024). Here we empirically find similar phenomena in certain deep ReLU networks. Analytical proof of alignment in deep ReLU networks is an intriguing future direction.
@@ 6 Discussion
Implication of Bias Removal. We studied the implications of removing bias in ReLU networks in terms of the expressivity and learning dynamics. Theorem 1 shows that two-layer bias-free (leaky) ReLU networks cannot express any odd functions except for linear functions. Theorem 7 shows that for datasets with an even input distribution and an odd target function, two-layer bias-free (leaky) ReLU networks have the same time evolution as a linear network (modulo scale factors) under initialization Assumption 5. We also presented examples in which the bias-free ReLU network evolves like multiple independent linear networks, in Section 4.2. In these cases, comparing a bias-free ReLU network with its linear counterpart provides an intuitive understanding of the behavior of ReLU networks. On the flip side, the simplicity of bias-free ReLU networks suggests that ReLU networks with bias may exhibit more complicated behaviors, which are not fully addressed by studies on bias-free networks, and remain open questions.
@@
One common argument in studies of bias-free ReLU networks is that we can stack the input ${\bm{x}}$ with an additional one, i.e., $[{\bm{x}},1]$. Then results derived for bias-free networks could extend to networks with bias and the removal of bias might thus have a minor implication. This argument is valid for some studies (Allen-Zhu et al., 2019; Zou et al., 2020), but not all. For example, Soudry et al. (2018) found that two-layer bias-free ReLU networks trained with logistic loss converge to the max-margin classifier on linearly separable datasets. As clarified by Soudry et al. (2018), this technical result holds when the inputs are stacked with an additional one. However, the max-margin solution for the dataset with stacked inputs is not the max-margin solution for the original dataset. Thus, the convergence to max-margin solution result does not directly extend to ReLU networks with bias.
Perturbed Symmetric Dataset. We have shown an exact equivalence between two-layer bias-free (leaky) ReLU networks and linear networks under symmetry Condition 3 on the dataset in Theorem 7. In practice, no datasets satisfies Condition 3 precisely. However, two-layer bias-free ReLU networks may still struggle to fit a dataset that approximately satisfies Condition 3. We present a simple example with six data points in Figure 5. The ReLU network loss curve closely follows the linear network loss curve in the early phase, when it first learns a nearly linear solution, as shown in LABEL:fig:asym_loss and LABEL:fig:asym_t10k. After a plateau, the ReLU network diverges from the linear network dynamics and converges to a nonlinear solution. The more symmetric the dataset, the longer the plateau a two-layer bias-free ReLU network undergoes before learning a nonlinear solution. As shown in LABEL:fig:asym_scaling, the inverse of the plateau duration scales approximately linearly with the deviation of the asymmetric data point, $\Delta y$. The scaling becomes less accurate for larger $\Delta y$ because the corresponding dataset more severely violates symmetric Condition 3, where the ReLU network no longer behaves like a linear network. When the dataset is exactly symmetric, the ReLU network never learns a nonlinear solution. Furthermore, as shown in LABEL:fig:asym_t100k, the decision boundaries at convergence are close to the data points, and thus probably not robust.
@@
Figure 5: Two-layer bias-free linear/ReLU network trained on a dataset that slightly violates the symmetry Condition 3. The $+$ and -- signs represent data points with $+1$ and -- labels respectively. The right middle data point is slight asymmetric with coordinates $(1,\Delta y)$. (a) Loss curves of the ReLU network with different $\Delta y$ and the linear network with $\Delta y=0.1$. (b) The duration of the plateau, during which the ReLU network implements a nearly linear solution, scales approximately with $1/\Delta y$. (c,d) The ReLU network output during and at the end of training. This ReLU network is trained on the dataset with $\Delta y=0.1$. Experimental details are provided in Appendix H.
@@ Appendix A Additional Related Work
Implicit / Simplicity Bias. Many works have studied the implicit bias or simplicity bias of two-layer bias-free ReLU networks under various assumptions on the dataset. Brutzkus et al. (2018); Wang et al. (2019); Lyu et al. (2021); Sarussi et al. (2021); Wang & Ma (2023) considered linearly separable binary classification tasks. Phuong & Lampert (2021); Wang & Pilanci (2022); Min et al. (2024) studied orthogonally separable classification (i.e., where for every pair of labeled examples $({\bm{x}}_{i},y_{i}),({\bm{x}}_{j},y_{j})$ we have ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}>0$ if $y_{i}=y_{j}$ and ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}\leq 0$ if otherwise). Boursier et al. (2022); Frei et al. (2023b; c); Kou et al. (2023) studied binary classification with exactly or nearly orthogonal input (i.e., where ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}=0$ if $i\neq j$). Orthogonal input is a sufficient condition for linear separability for binary classification tasks. Frei et al. (2023a); Meng et al. (2024); Xu et al. (2024) studied XOR-like datasets. Vardi et al. (2022); Frei et al. (2023d) studied datasets with adversarial noise. We add to this line of research by studying a case with extreme simplicity bias, i.e., behaving like linear networks.
Low-Rank Weights. Maennel et al. (2018) is the seminal work on the low-rank weights in two-layer ReLU networks trained from small initialization. They described the phenomenon as “quantizing", where the first layer weight vectors align with a small number of directions in the early phase of training. Luo et al. (2021) identified when two-layer bias-free ReLU networks form low-rank weights in terms of the initialization and the network width. Timor et al. (2023) provided cases where gradient flow on two-layer and deep ReLU networks provably minimize or not minimize the ranks of weight matrices. Frei et al. (2023c); Kou et al. (2023) computed the numerical rank of the converged weights of two-layer bias-free ReLU networks for nearly orthogonal datasets, and found that weights in leaky ReLU networks have rank at most two and weights in ReLU networks have a numerical rank upper bounded by a constant. Chistikov et al. (2023) showed two-layer bias-free ReLU networks are implicitly biased to learn the network of minimal rank under the assumption that training points are correlated with the teacher neuron. Min et al. (2024); Boursier & Flammarion (2024) studied the early phase learning dynamics to understand how the low-rank weights form. Petrini et al. (2022; 2023) conducted experiments on practical datasets to show that two-layer bias-free ReLU networks learn sparse features, which can be detrimental and lead to overfitting. Le & Jegelka (2022) generalize the low-rank phenomenon in linear and ReLU networks to arbitrary non-homogeneous networks whose last few layers contain linear fully-connected and linear ResNet blocks.

### v3

@@ 5 Learning Dynamics in Deep Bias-Free ReLU Networks
In Section 4, we showed that two-layer bias-free ReLU networks behave like linear networks under symmetry Condition 3 and small initialization. This does not extend to deep bias-free networks. When trained on a dataset satisfying Condition 3, deep bias-free ReLU networks can learn nonlinear solutions if the target function is nonlinear, as shown in LABEL:fig:express-fan (upper right). Nonetheless, we find deep bias-free ReLU networks can form low-rank weights that are similar to those in deep linear networks. We give an example where the empirical input distribution is even and the target function is linear.
@@
where $u>0$ represents the norm of each layer, and ${\bm{r}},{\bm{r}}_{1},{\bm{r}}_{2},\cdots,{\bm{r}}_{L}$ are unit norm column vectors. The vectors ${\bm{r}}_{l}^{+},{\bm{r}}_{l}^{-}$ denote the positive and negative elements in ${\bm{r}}_{l}$. The equal norm $u$ of all layers is a consequence of small initialization (Du et al., 2018). Note that the weights can be written in blocks, as Equation 17, only after permuting the positive and negative elements. We use this permuted notation for the sake of exposition; no additional assumptions are required.
In a deep ReLU network, we empirically find that when the weights are trained from small initialization and the target function is linear, the weights exhibit a particular rank-one and rank-two structure, as shown in Figures 5 and 10. For the first and last layers, the weights in the deep ReLU network have the same rank-one structure as their linear counterpart. For the intermediate layers, weights in the deep ReLU network are rank-two. Specifically, positive weights in the ReLU network correspond to positive weights in the linear network and zero weights in the ReLU network correspond to negative weights in the linear network. Based on the empirical observation, we propose the following conjecture on the weights of the deep ReLU network.
Conjecture 11.
A deep bias-free ReLU network is trained from small initial weights on a dataset where the empirical input distribution is even, $p({\bm{x}})=p(-{\bm{x}})$, and the target function is linear, that is the target output is generated as $y={{\bm{w}}^{*}}^{\top}{\bm{x}}$. We conjecture that the weights at a certain time $t_{0}$ during training take the following form:
@@
${\bm{W}}_{L}$ $=u{\bm{r}}_{L-1}^{\top}=u\begin{bmatrix}{{\bm{r}}_{L-1}^{+}}^{\top}&{{\bm{r}}_{L-1}^{-}}^{\top}\end{bmatrix},$ (18c)
where the notations are consistent with Equation 17. We also conjecture that $\|{\bm{r}}_{l}^{+}\|=\|{\bm{r}}_{l}^{-}\|,\,l=1,2,\cdots,L-1$.
Proposition 12.
If Conjecture 11 is true, then for all $t\geq t_{0}$, the weights of the deep bias-free ReLU network will maintain the form in Equation 18 and the network implements a linear function given by
@@
Moreover, its learning dynamics is described by
$\tau\dot{\bm{W}}_{l}=\left(\prod_{l^{\prime}=l+1}^{L}{\bm{W}}_{l^{\prime}}\right)^{\top}\left(\frac{1}{2}{\bm{\beta}}^{\top}-\frac{1}{4}\prod_{l^{\prime}=1}^{L}{\bm{W}}_{l^{\prime}}{\bm{\Sigma}}\right)\left(\prod_{l^{\prime}=l-1}^{L}{\bm{W}}_{l^{\prime}}\right)^{\top},$ (20)
which is the same as that of a deep linear network in Equation 7, modulo the constant coefficients.
@@
We provide the proof of Proposition 12 in Appendix E and offer a brief explanation for Equation 19 here. In the first equality of Equation 19, we dropped the activation functions except for the one between the first and second layers. This is because the second layer weights, ${\bm{W}}_{2}$, is non-negative as shown in LABEL:fig:deep-relu-rank12, and so is the output of a ReLU activation function, $\sigma({\bm{W}}_{1}{\bm{x}})$. Hence, their product, ${\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})$, is also non-negative. We thus have $\sigma({\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}}))={\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})$. The same applies to all subsequent layers. The second equality is obtained by substituting the weights defined in Equation 18 into the expression.
Figure 5: Low-rank weights in deep linear and ReLU bias-free networks. A three-layer linear network and a three-layer ReLU network are trained on the same dataset starting from the same small random weights. The dataset has a linear target function and an even empirical input data distribution. We plot the weights when the loss has approached zero. ${\bm{W}}_{1}$, ${\bm{W}}_{3}$, and positive elements in ${\bm{W}}_{2}$ have approximately the same structure in the linear and ReLU networks. Elements of ${\bm{W}}_{2}$ that are negative in the linear network are approximately zero in the ReLU network. The neurons are permuted for visualization. Simulations with deeper networks are presented in Figure 10. Experimental details are provided in Appendix H.
@@
While the proof of Conjecture 11 remains an open question, we provide some intuition to interpret it. Specifically, we notice that in deep linear networks and certain deep ReLU networks, the weights of an intermediate layer align with the inputs to that layer. For example, the second-layer weight in a deep linear network is ${\bm{W}}_{2}^{\mathrm{lin}}=u{\bm{r}}_{2}{\bm{r}}_{1}^{\top}$ as given in Equation 17. Every row of ${\bm{W}}_{2}^{\mathrm{lin}}$ aligns with ${\bm{r}}_{1}^{\top}$, which is parallel to any input to the second layer, ${\bm{W}}_{1}^{\mathrm{lin}}{\bm{x}}=u{\bm{r}}_{1}{\bm{r}}^{\top}{\bm{x}}$. The second-layer weight in the deep ReLU network is given in Equation 18. Some rows of ${\bm{W}}_{2}$ align with $\begin{bmatrix}{{\bm{r}}_{1}^{+}}^{\top}&{\bm{0}}\end{bmatrix}$, which is parallel to some inputs (${\bm{r}}^{\top}{\bm{x}}>0$) to the second layer, $\sigma({\bm{W}}_{1}{\bm{x}})=u\begin{bmatrix}{\bm{r}}_{1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}$. Other rows of ${\bm{W}}_{2}$ align with $\begin{bmatrix}{\bm{0}}&{{\bm{r}}_{1}^{-}}^{\top}\end{bmatrix}$, which is parallel with inputs (${\bm{r}}^{\top}{\bm{x}}<0$) to the second layer, $\sigma({\bm{W}}_{1}{\bm{x}})=u\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}$. This alignment phenomenon has been proven for deep linear networks (Ji & Telgarsky, 2019; Marion & Chizat, 2024), and we here find similar phenomena empirically in deep ReLU networks.
@@ 6.1 Implication of Bias Removal
We studied the implications of removing bias in ReLU networks in terms of the expressivity and learning dynamics. Theorem 1 shows that two-layer bias-free (leaky) ReLU networks cannot express any odd functions except for linear functions. Theorem 8 shows that for datasets with an even input distribution and an odd target function, two-layer bias-free (leaky) ReLU networks have the same time evolution as a linear network (modulo scale factors) under initialization Assumption 7. We also presented examples in which the bias-free ReLU network evolves like multiple independent linear networks, in Section 4.2. In these cases, comparing a bias-free ReLU network with its linear counterpart provides an intuitive understanding of the behavior of ReLU networks. On the flip side, the simplicity of bias-free ReLU networks suggests that ReLU networks with bias may exhibit more complicated behaviors, which are not fully addressed by studies on bias-free networks, and remain open questions.
@@
One common argument in studies of bias-free ReLU networks is that we can stack the input ${\bm{x}}$ with an additional one, i.e., $\tilde{\bm{x}}=[{\bm{x}},1]$. The behaviors of a biased ReLU network with input ${\bm{x}}$ are the same as those of a bias-free ReLU network with input $\tilde{\bm{x}}$, suggesting that the implication of bias removal might be trivial. While this argument is valid in certain settings (Allen-Zhu et al., 2019; Zou et al., 2020), it comes with important caveats in others. For example, Soudry et al. (2018) showed that two-layer bias-free ReLU networks trained with logistic loss converge to the max-margin classifier on linearly separable datasets. As clarified by Soudry et al. (2018), this technical result holds when the inputs are stacked with an additional one. However, the max-margin solution of the dataset $\{\tilde{\bm{x}}_{\mu},y_{\mu}\}_{\mu=1}^{P}$ is not the max-margin solution of the original dataset $\{{\bm{x}}_{\mu},y_{\mu}\}_{\mu=1}^{P}$. Hence, a ReLU network with bias converges to a solution different from the max-margin solution obtained by the bias-free ReLU network. This distinction highlights that the presence of bias terms can change the inductive bias of the ReLU network, leading to convergence to different solutions.
6.2 Perturbed Symmetric Dataset
@@
Figure 6: Two-layer bias-free linear/ReLU network trained on a dataset that slightly violates the symmetry Condition 3. The $+$ and -- signs represent data points with $+1$ and -- labels respectively. The right middle data point is slight asymmetric with input coordinates $(1,\delta)$. (a) Loss curves of the ReLU network with different $\delta$ and the linear network with $\delta=0.1$. (b) The duration of the plateau, during which the ReLU network implements a nearly linear solution, scales approximately with $1/\delta$. (c,d) The ReLU network output during and at the end of training. This ReLU network is trained on the dataset with $\delta=0.1$. Experimental details are provided in Appendix H.
We have shown an exact equivalence between two-layer bias-free (leaky) ReLU networks and linear networks under symmetry Condition 3 on the dataset in Theorem 8. In practice, no datasets satisfies Condition 3 precisely. However, two-layer bias-free ReLU networks may still struggle to fit a dataset that approximately satisfies Condition 3. We present a simple example with six data points in Figure 6. The ReLU network loss curve closely follows the linear network loss curve in the early phase, when it first learns a nearly linear solution, as shown in LABEL:fig:asym_loss and LABEL:fig:asym_t10k. After a plateau, the ReLU network diverges from the linear network dynamics and converges to a nonlinear solution. The more symmetric the dataset, the longer the plateau a two-layer bias-free ReLU network undergoes before learning a nonlinear solution. When the dataset is exactly symmetric, the ReLU network never learns a nonlinear solution. As shown in LABEL:fig:asym_scaling, the inverse of the plateau duration scales approximately linearly with the deviation of the asymmetric data point, $\delta$. The scaling becomes less accurate for larger $\delta$ because the corresponding dataset more severely violates symmetric Condition 3, where the ReLU network no longer behaves like a linear network. Furthermore, as shown in LABEL:fig:asym_t100k, the decision boundaries at convergence are close to the data points, and thus probably not robust.
We note a concurrent work (Boursier & Flammarion, 2024b) that considers two-layer bias-free ReLU networks trained on symmetric datasets with a different form of perturbation. They assume the empirical input distribution is even and the target output is generated as $y={{\bm{w}}^{*}}^{\top}{\bm{x}}+\xi$, where $\xi$ is the noise sampled independently from a zero-mean Gaussian distribution. In their setup, the ReLU network converges to the ordinary least square linear estimator when the number of training samples exceeds a certain threshold. In their case, the mean of the noise approaches zero as the number of training samples increases, which reduces the asymmetric perturbation and leads to convergence to a linear solution. By analogy, in our case, a smaller $\delta$ means a smaller asymmetric perturbation, which leads to a longer plateau during which the ReLU network is stuck at a linear solution.
Acknowledgement
The authors are grateful to Peter Orbanz, Ingo Steinwart, Spencer Frei, Rodrigo Carrasco-Davis, Lukas Braun, Clémentine Dominé, Zheng He, and William Dorrell for helpful discussions. The authors thank the following funding sources: Gatsby Charitable Foundation (GAT3850) to YZ, AS, and PEL; Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) to AS; Schmidt Science Polymath Award to AS; Wellcome Trust (110114/Z/15/Z) to PEL.
@@ Appendix A Additional Related Work
Implicit / Simplicity Bias. Many works have studied the implicit bias or simplicity bias of two-layer bias-free ReLU networks under various assumptions on the dataset. Brutzkus et al. (2018); Wang et al. (2019); Lyu et al. (2021); Sarussi et al. (2021); Wang & Ma (2023) considered linearly separable binary classification tasks. Phuong & Lampert (2021); Wang & Pilanci (2022); Min et al. (2024) studied orthogonally separable classification (i.e., where for every pair of labeled examples $({\bm{x}}_{i},y_{i}),({\bm{x}}_{j},y_{j})$ we have ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}>0$ if $y_{i}=y_{j}$ and ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}\leq 0$ if otherwise). Boursier et al. (2022); Frei et al. (2023b; c); Kou et al. (2023); Dana et al. (2025) studied binary classification with exactly or nearly orthogonal input (i.e., where ${\bm{x}}_{i}^{\top}{\bm{x}}_{j}=0$ if $i\neq j$). Orthogonal input is a sufficient condition for linear separability for binary classification tasks. Refinetti et al. (2021); Frei et al. (2023a); Meng et al. (2024); Xu et al. (2024) studied XOR-like datasets. Vardi et al. (2022); Frei et al. (2023d) studied datasets with adversarial noise. We add to this line of research by studying a case with extreme simplicity bias, i.e., behaving like linear networks.
Low-Rank Weights. Maennel et al. (2018) is the seminal work on the low-rank weights in two-layer ReLU networks trained from small initialization. They described the phenomenon as “quantizing", where the first layer weight vectors align with a small number of directions in the early phase of training. Luo et al. (2021) identified when two-layer bias-free ReLU networks form low-rank weights in terms of the initialization and the network width. Timor et al. (2023) provided cases where gradient flow on two-layer and deep ReLU networks provably minimize or not minimize the ranks of weight matrices. Frei et al. (2023c); Kou et al. (2023) computed the numerical rank of the converged weights of two-layer bias-free ReLU networks for nearly orthogonal datasets, and found that weights in leaky ReLU networks have rank at most two and weights in ReLU networks have a numerical rank upper bounded by a constant. Chistikov et al. (2023) showed two-layer bias-free ReLU networks are implicitly biased to learn the network of minimal rank under the assumption that training points are correlated with the teacher neuron. Min et al. (2024); Boursier & Flammarion (2024a) studied the early phase learning dynamics to understand how the low-rank weights form. Petrini et al. (2022; 2023) conducted experiments on practical datasets to show that two-layer bias-free ReLU networks learn sparse features, which can be detrimental and lead to overfitting. Le & Jegelka (2022) generalize the low-rank phenomenon in linear and ReLU networks to arbitrary non-homogeneous networks whose last few layers contain linear fully-connected and linear ResNet blocks.